OpenStudy (anonymous):

PLEASE HELP ME HOW TO DO THIS, DONT JUST GIVE ME THE ANSWER. THE QUESTION IS GOING TO BE DRAWN

OpenStudy (anonymous):

subtract the exponents

OpenStudy (anonymous):

\[\frac{\sqrt[9]{x^7}}{\sqrt[9]{x^5}}=\sqrt[9]{\frac{x^7}{x^5}}\]

OpenStudy (anonymous):

the "0" should be a "9"

OpenStudy (anonymous):

\[\frac{x^7}{x^5}=x^{7-5}=x^2\] is inside the radical, but the radical stays the same

OpenStudy (anonymous):

ohhh okay. thank you. can you hellp me ith a couple more please? :)

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):